OpenStudy (anonymous):

solve for x. 5x^2=-13x-6

OpenStudy (anonymous):

x = - 3/5 and x = -2

OpenStudy (anonymous):

5x² = -13x - 6 Getting all the terms to one side. 0 = 5x² + 13x + 6 Using the tic tac toe method to factor. See < http://www.slideshare.net/qvamath/tic-tac-toe-factoring >. 5x²| 6|30x² 5x| 2|10x x| 3|3x 0 = (5x + 3)(x + 2) Zero product property tells us that if two or more things multiply to 0, then one of them must equal zero. 5x + 3 = 0 5x = -3 x = (-3/5) x + 2 = 0 x = -2 x = -2, (-3/5)